# On the Design of One-step Diffusion via Shortcutting Flow Paths

### v3

@@ 1 Introduction
Diffusion-based models have become the dominant paradigm in deep generative modeling (sohldickstein2015deepunsupervisedlearningusing; hodiffusion; sde), progressively transforming samples from a prior distribution toward the data distribution. However, dozens or even hundreds of neural function evaluations (NFEs) are typically required, resulting in slow inference and limited real-time use (song2020generativemodelingestimatinggradients; salimans2022progressivedistillationfastsampling; dpmplus; dpmv3). Consistency models (cm; icm) are pioneering works that attempt to achieve one-step generation (Luo2023LatentCM; wang2023diffusiongantraininggansdiffusion; yin2024improved; yin2024onestepdiffusiondistributionmatching; salimans2024multistepdistillationdiffusionmodels; geng2023onestep; geng2025consistency), but a costly two-stage training process is required, i.e., first training a reliable diffusion model and then distilling velocity or score from it. Despite the costly two-stage training, they offer fast generation, which motivates further research into improving training efficiency.
Recently, one-step diffusion models trained from scratch have emerged, such as Consistency Training (CT) (cm) as the training-from-scratch variant of consistency models, Inductive Moment Matching (IMM) (imm), and Shortcut Diffusion (SCD) (scd). These models aim to learn direct shortcut mappings between intermediate states along the probability flow trajectories of the probability flow, thus enabling one-step generation; we refer to such models as shortcut models. Building on this principle, continuous-time shortcut models such as sCT (scm) and MeanFlow (meanflow) have been introduced, achieving state-of-the-art performance in one-step generation for image synthesis. Their efficiency and effectiveness in both training and generation have stimulated further exploration in improving their sampling fidelity.
@@ 2.1 Shortcutting flows with flow map solvers
Let $p_{\text{data}}(\bm{x})$ be the data distribution, and $p_{\text{prior}}=\mathcal{N}(\bm{0},\sigma^{2}\bm{\mathrm{I}})$ be a Gaussian distribution with zero mean and variance $\sigma^{2}$. In the following, we write $\sigma=1$ by default for notational simplicity. According to stochastic interpolants (stochasticinterp), diffusion models establish a probabilistic path between $p_{0}=p_{\text{data}}$ and $p_{1}=p_{\text{prior}}$ such that $\bm{x}_{t}=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon},$ where $\bm{x}_{0}\sim p_{0}$, $\bm{\varepsilon}\sim p_{1}$, and $\alpha_{t},\sigma_{t}\geq 0$; with boundary conditions $\alpha_{0}=\sigma_{1}=1$ and $\alpha_{1}=\sigma_{0}=0$. Both the forward noising and inverse denoising processes are governed by the probability flow ODE (PF-ODE) as $\dot{\bm{x}}_{t}=\bm{v}_{t}(\bm{x}_{t})$, where $\bm{v}_{t}(\bm{x})$ is the marginal velocity $\bm{v}_{t}(\bm{x})=\dot{\alpha}_{t}\mathbb{E}(\bm{x}_{0}|\bm{x}_{t}=\bm{x})+\dot{\sigma}_{t}\mathbb{E}(\bm{\varepsilon}|\bm{x}_{t}=\bm{x})$.
@@
Probabilistic paths satisfying the above are defined as flow paths. For example, with the reformulation by scm, the EDM preconditioner path (edm) can be transformed to a cosine path (sit) with $\sigma=\sigma_{\text{data}}$, $\alpha_{t}=\cos(\frac{\pi}{2}t)$, and $\sigma_{t}=\sin(\frac{\pi}{2}t)$; in Rectified Flow (recflow), $\sigma=1$, $\alpha_{t}=1-t$ and $\sigma_{t}=t$, leading to the linear path (fm; ifm). Since $\bm{v}_{t}(\bm{x})$ is inaccessible, the conditional path is established for tractable training, where the corresponding conditional velocity is $\bm{v}_{t|0}=\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0})=\dot{\alpha}_{t}\bm{x}_{0}+\dot{\sigma}_{t}\bm{\varepsilon}$ that neural networks $F^{\theta}(\bm{x}_{t},t)$ are trained to approximate. In sampling, one can first sample $\bm{x}_{1}=\bm{\varepsilon}\sim p_{\text{prior}}$, and then simulate a trajectory of the flow through the PF-ODE as $\dot{\bm{x}}_{t}=F^{\theta}(\bm{x}_{t},t)$.
@@
In order to shortcut established flow paths from time $t$ to $r$ ($0\leq r\leq t\leq 1$), we introduce the flow map notation (flowmap; tutorial; boffi2025how) to express the design frame for simplicity. A flow map $X_{t,r}$ is defined as the unique map such that $X_{t,r}(\bm{x}_{t})=\bm{x}_{r}$, for all $(t,r)\in[0,1]^{2}$, where $\bm{x}_{r}$ is the solution of PF-ODE, which corresponds to position in physics. According to the PF-ODE, one can easily derive the flow map solution through
@@
With the definition of average velocity over time (meanflow) as $\bm{u}_{t,r}$, we can rewrite Eq. 1 to express the flow map solution $\bm{x}_{r}={X}_{t,r}(\bm{x}_{t})$ through
@@
or to infer ${X}_{t,r}(\bm{x}_{t})$ with the instantaneous velocity $\bm{v}_{t}$, through DDIM-solver (ddim) as first-order approximation of DPM-solver (dpm), which reads
@@ 2.2 Learning to shortcut flow paths
We claim that the previous methods shortcut the flow paths of a diffusion model by regularizing a one-step flow map prediction against the two-step flow map target. In practice, they first sample time points $r,s,t\sim p(\tau)$ with $r\leq s\leq t$, and then use the consistency property (tutorial; flowmap) as detailed in Appendix A.1 to design a shortcut model trained from scratch, which reads
@@
We denote by $F^{\theta}$ the neural network with parameters $\theta$, whose architecture is instantiated as U-Net (sde) in the pixel space, or as DiT (dit) / SiT (sit) in the latent space. To obtain the flow map, Eq. 1, Eq. 2, and Eq. 3 can all serve as solutions. Since the integral term $\int_{t}^{r}\bm{v}_{\tau}(\bm{x}_{\tau})d\tau$ in Eq. 1 is intractable in general, the DDIM solver with instantaneous velocity is adopted practically when estimating the flow map with $\bm{v}_{t}^{\theta}$ parameterized by $F^{\theta}$ or the conditional velocity $\bm{v}_{t|0}$ through Eq. 3. Alternatively, if $F^{\theta}$ parameterizes average velocity $\bm{u}^{\theta}_{t,r}$, the flow map can be obtained directly through Eq. 2.
@@ 2.3 Examples: discrete- and continuous-time shortcut models
where the loss metric is LPIPS (lpips) applied directly in pixel space with $w=1$, and $l_{\mathrm{ct}}(\bm{x}_{t},r,s,t;\theta)$ coincides with its formulation in CTs with details shown in Appendix B.1. Alternatively, if we parameterize the average velocity as $\bm{u}^{\theta}_{t,r}(\bm{x}_{t})=F^{\theta}(\bm{x}_{t},t,r)$, and estimate both the one-step prediction and the two steps in the target with neural networks $F^{\theta}$, we thus obtain the $l(\bm{x}_{t},r,s,t;\theta)$ in SCD through Eq. 2. Due to equi-spacing time points as ts=sr=ht-s=s-r=h, it reads
@@
where the loss metric is set as squared $l_{2}$-distance, $w=\frac{1}{4h^{2}}$, and $\hat{\bm{x}}_{s}=\bm{x}_{t}+(s-t)\bm{u}^{\theta}_{s,t}(\bm{x}_{t})$, with details shown in Appendix B.2. Moreover, for IMMs, conditional samples $\{\bm{x}^{(i)}_{0},\bm{\varepsilon}^{(i)}\}_{i=1}^{B}$ within a mini-batch of size $B$ are first partitioned into different groups, and $\{r,s,t\}$ are drawn for each group. With similar flow map construction to CTs, the loss metric is implemented with a grouped kernel function as MMD (mmd), applied to measure both inter- and intra-sample similarities between $\{\bm{\hat{x}}_{r},\bm{x}^{\theta}_{r}\}$ within the group, as further detailed in Appendix B.3.
@@ 2.4 Discussion: shortcutting flow paths under marginal velocity fields
However, such pairs $\{\bm{x}_{t},\bm{x}_{r}\}$ cannot be obtained via simulation-based sampling: once $\bm{x}_{t}$ is sampled, $\bm{x}_{r}$ remains inaccessible because both $\bm{v}_{t}(\bm{x})$ and its integral from $t$ to $r$ are intractable. To overcome this, a common design paradigm is employed, which is to let the network’s outputs, or the conditional velocity alternatively, estimate $\bm{x}_{r}$ in two steps: first producing an intermediate $\hat{\bm{x}}_{s}$, and then constructing an estimated target $\hat{\bm{x}}_{r}$, as shown in Eq. 5 in Sec. 2.2. This makes training feasible. Although one may also construct multi-step (i.e., more than two) flow map targets for simulating the $\{\bm{x}_{t},\bm{x}_{r}\}$ pairs (kim2023consistency), the paradigms of two-step target construction approximated by one-step prediction are sufficiently general according to the following theoretical justification with detailed proof in Appendix C.2. Note that we classify the aforementioned methods’ training objective into DTSC and CTSC. For example, $l_{\text{mf}}$ and $l_{\text{sct}}$ are instances of $l_{\text{ctsc}}$.
@@
From another perspective, this explains why distilling from a pretrained diffusion model is often more effective than training from scratch (cm; scm). Unlike (s)CT, which are trained from scratch, (s)CM benefits from distillation by learning from a pretrained velocity field $\bm{v}_{t}^{\phi}(\bm{x})$. In practical training, the conditional velocity $\bm{v}_{t|0}$ and network output $\bm{v}_{t}^{\theta}$ in $\mathrm{sg}(\cdot)$ in Eq. 6 and 9 are replaced with $\bm{v}_{t}^{\phi}$, which closely approximates $\bm{v}_{t}(\bm{x}_{t})$. This substantially reduces errors in estimating the two-step flow targets, providing more accurate supervision for network training.
@@ 3 Elucidating the design space of shortcut models
Empirically, we evaluate the proposed formulation using a unified codebase implementation with the same training iterations and batch sizes. For unconditional generation on CIFAR-10, we employ U-Nets (sde) ($\sim$55M param.) as the network architecture operating directly in the pixel space. For conditional generation in ImageNet-$\times256$, with and without classifier-free guidance, we use a SiT-B/2 (sit) architecture ($\sim$131M param.), operating in the latent space via a pretrained VQVAE (stablediff). While sCT is originally initialized from the teacher diffusion model as stated in scm, we train all the discussed models from scratch, for a fair comparison. Fig. 4(a) summarizes the results of one-step generation on the two datasets, with additional setting of classifier-guidance-free learning, as discussed in meanflow. Further details on settings are provided in Appendix D.1.
@@
Linear paths are generally regarded as more analytically tractable and easier to employ for training and sampling tricks (e.g., classifier-free guidance), owing to their simple formulation. By contrast, in pixel-space generative modeling, cosine paths are often considered more stable for training convergence, because they induce a stochastic process with fixed variance. Exploration of these two flow paths in the context of shortcut models remains underexplored. Here we extend cosine-path-based models (CT and sCT) to their linear-path counterparts. Fig. 4(a) shows that shortcut models with linear paths are more competitive. We attribute this to the fact that the marginal velocity fields generated by linear paths as conditional paths are at lower convex transport cost (recflow), implying lower curvature of the velocity-field-governed trajectories. Consequently, the simulated two-step flow map targets are less likely to deviate from the ideal. Furthermore, while cosine paths are optimal in the setting of diffusion and flow matching (santos2023usingornsteinuhlenbeckprocessunderstand) under Fisher information metrics, we theoretically justify that linear paths in the setting of shortcut models are optimal conditioned on data samples in Appendix C.3. Based on this, our subsequent analysis will focus on the linear path.
@@
Under mild regularity conditions shown in Appendix C.4.1, the Wasserstein-2 distance of shortcut models with one-step generation is bounded as: $W_{2}^{2}(p_{0},p^{\theta}_{0})\leq 2\left(\mathrm{BV_{\text{ctsc}}}+8\mathrm{Var}\big[\frac{d}{dt}\bm{u}_{t,r}^{\theta}(\bm{x}_{t})\big]+8\sigma_{\bm{v}_{t|0}}^{2}\right)\Big|_{r=0,t=1},$ (10) $W_{2}^{2}(p_{0},p^{\theta}_{0})\leq 2\Big(\mathrm{BV_{\text{dtsc}}}+8\delta_{2}^{2}\,\mathrm{Var}\big[\bm{u}_{s,r}^{\theta}(\bm{x}_{t})\big]+8(1+\ell^{2}\delta_{2}^{2})\delta_{1}^{2}\,\sigma_{\text{dtsc}}^{2}\Big)\Big|_{r=0,t=1},$~ (11) where $\mathrm{BV_{\bullet}}=\mathrm{Bias}^{2}_{\bullet\text{-tgt}}+\mathrm{Bias}^{2}_{\bullet\text{-loss}}+2\mathrm{Var}\!\big[\bm{u}_{t,r}^{\theta}(\bm{x}_{1})\big]$ with $\bullet\in\{\mathrm{ctsc},\mathrm{dtsc}\}$, and $\mathrm{Bias}^{2}_{\bullet\text{-tgt}}$ and $\mathrm{Bias}^{2}_{\bullet\text{-loss}}$ are defined in Prop.C.8 ; $\delta_{1}=t-s$, $\delta_{2}=s-r$; $\ell$ is the local Lipschitz constant of $\bm{u}^{\theta}$; $\sigma_{\bm{v}_{t|0}}^{2}$ is the variance of the conditional velocity, defined by $\sigma_{\bm{v}_{t|0}}^{2}\coloneqq\mathrm{Var}(\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0}))$; $\sigma_{\text{dtsc}}^{2}=\sigma_{\bm{v}_{t|0}}^{2}$ when using CT’s two-step flow map targets, or $\sigma_{\text{dtsc}}^{2}=\mathrm{Var}\!\big[\bm{u}_{t,s}^{\theta}(\bm{x}_{t})\big]$ when using SCD’s flow map targets.

### v5

@@ 1 Introduction
Diffusion-based models have become the dominant paradigm in deep generative modeling (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), progressively transforming samples from a prior distribution toward the data distribution. However, dozens or even hundreds of neural function evaluations (NFEs) are typically required, resulting in slow inference and limited real-time use (Song and Ermon, 2020; Salimans and Ho, 2022; Lu et al., 2025; Zheng et al., 2023). Consistency models (Song et al., 2023; Song and Dhariwal, 2023) are pioneering works that attempt to achieve one-step generation (Luo et al., 2023; Wang et al., 2023; Yin et al., 2024a; c; Salimans et al., 2024; Geng et al., 2023; 2025b), but a costly two-stage training process is required, i.e., first training a reliable diffusion model and then distilling velocity or score from it. Despite the costly two-stage training, they offer fast generation, which motivates further research into improving training efficiency.
Recently, one-step diffusion models trained from scratch have emerged, such as Consistency Training (CT) (Song et al., 2023) as the training-from-scratch variant of consistency models, Inductive Moment Matching (IMM) (Zhou et al., 2025), and Shortcut Diffusion (SCD) (Frans et al., 2025). These models aim to learn direct shortcut mappings between intermediate states along the probability flow trajectories of the probability flow, thus enabling one-step generation; we refer to such models as shortcut models. Building on this principle, continuous-time shortcut models such as sCT (Lu and Song, 2025) and MeanFlow (Geng et al., 2025a) have been introduced, achieving state-of-the-art performance in one-step generation for image synthesis. Their efficiency and effectiveness in both training and generation have stimulated further exploration in improving their sampling fidelity.
@@ 2.1 Shortcutting flows with flow map solvers
Let $p_{\text{data}}(\bm{x})$ be the data distribution, and $p_{\text{prior}}=\mathcal{N}(\bm{0},\sigma^{2}\bm{\mathrm{I}})$ be a Gaussian distribution with zero mean and variance $\sigma^{2}$. In the following, we write $\sigma=1$ by default for notational simplicity. According to stochastic interpolants (Albergo et al., 2023), diffusion models establish a probabilistic path between $p_{0}=p_{\text{data}}$ and $p_{1}=p_{\text{prior}}$ such that $\bm{x}_{t}=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon},$ where $\bm{x}_{0}\sim p_{0}$, $\bm{\varepsilon}\sim p_{1}$, and $\alpha_{t},\sigma_{t}\geq 0$; with boundary conditions $\alpha_{0}=\sigma_{1}=1$ and $\alpha_{1}=\sigma_{0}=0$. Both the forward noising and inverse denoising processes are governed by the probability flow ODE (PF-ODE) as $\dot{\bm{x}}_{t}=\bm{v}_{t}(\bm{x}_{t})$, where $\bm{v}_{t}(\bm{x})$ is the marginal velocity $\bm{v}_{t}(\bm{x})=\dot{\alpha}_{t}\mathbb{E}(\bm{x}_{0}|\bm{x}_{t}=\bm{x})+\dot{\sigma}_{t}\mathbb{E}(\bm{\varepsilon}|\bm{x}_{t}=\bm{x})$.
@@
Probabilistic paths satisfying the above are defined as flow paths. For example, with the reformulation by Lu and Song (2025), the EDM preconditioner path (Karras et al., 2022) can be transformed to a cosine path (Ma et al., 2024) with $\sigma=\sigma_{\text{data}}$, $\alpha_{t}=\cos(\frac{\pi}{2}t)$, and $\sigma_{t}=\sin(\frac{\pi}{2}t)$; in Rectified Flow (Liu et al., 2022), $\sigma=1$, $\alpha_{t}=1-t$ and $\sigma_{t}=t$, leading to the linear path (Lipman et al., 2023; Tong et al., 2024). Since $\bm{v}_{t}(\bm{x})$ is inaccessible, the conditional path is established for tractable training, where the corresponding conditional velocity is $\bm{v}_{t|0}=\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0})=\dot{\alpha}_{t}\bm{x}_{0}+\dot{\sigma}_{t}\bm{\varepsilon}$ that neural networks $F^{\theta}(\bm{x}_{t},t)$ are trained to approximate. In sampling, one can first sample $\bm{x}_{1}=\bm{\varepsilon}\sim p_{\text{prior}}$, and then simulate a trajectory of the flow through the PF-ODE as $\dot{\bm{x}}_{t}=F^{\theta}(\bm{x}_{t},t)$.
@@
In order to shortcut established flow paths from time $t$ to $r$ ($0\leq r\leq t\leq 1$), we introduce the flow map notation (Boffi et al., 2025a; Liu, 2025; Boffi et al., 2025b) to express the design frame for simplicity. A flow map $X_{t,r}$ is defined as the unique map such that $X_{t,r}(\bm{x}_{t})=\bm{x}_{r}$, for all $(t,r)\in[0,1]^{2}$, where $\bm{x}_{r}$ is the solution of PF-ODE, which corresponds to position in physics. According to the PF-ODE, one can easily derive the flow map solution through
@@
With the definition of average velocity over time (Geng et al., 2025a) as $\bm{u}_{t,r}$, we can rewrite Eq. 1 to express the flow map solution $\bm{x}_{r}={X}_{t,r}(\bm{x}_{t})$ through
@@
or to infer ${X}_{t,r}(\bm{x}_{t})$ with the instantaneous velocity $\bm{v}_{t}$, through DDIM-solver (Song et al., 2021) as first-order approximation of DPM-solver (Lu et al., 2022), which reads
@@ 2.2 Learning to shortcut flow paths
We claim that the previous methods shortcut the flow paths of a diffusion model by regularizing a one-step flow map prediction against the two-step flow map target. In practice, they first sample time points $r,s,t\sim p(\tau)$ with $r\leq s\leq t$, and then use the consistency property (Liu, 2025; Boffi et al., 2025a) as detailed in Appendix A.1 to design a shortcut model trained from scratch, which reads
@@
We denote by $F^{\theta}$ the neural network with parameters $\theta$, whose architecture is instantiated as U-Net (Song et al., 2020) in the pixel space, or as DiT (Peebles and Xie, 2022) / SiT (Ma et al., 2024) in the latent space. To obtain the flow map, Eq. 1, Eq. 2, and Eq. 3 can all serve as solutions. Since the integral term $\int_{t}^{r}\bm{v}_{\tau}(\bm{x}_{\tau})d\tau$ in Eq. 1 is intractable in general, the DDIM solver with instantaneous velocity is adopted practically when estimating the flow map with $\bm{v}_{t}^{\theta}$ parameterized by $F^{\theta}$ or the conditional velocity $\bm{v}_{t|0}$ through Eq. 3. Alternatively, if $F^{\theta}$ parameterizes average velocity $\bm{u}^{\theta}_{t,r}$, the flow map can be obtained directly through Eq. 2.
@@ 2.3 Examples: discrete- and continuous-time shortcut models
where the loss metric is LPIPS (Zhang et al., 2018) applied directly in pixel space with $w=1$, and $l_{\mathrm{ct}}(\bm{x}_{t},r,s,t;\theta)$ coincides with its formulation in CTs with details shown in Appendix B.1. Alternatively, if we parameterize the average velocity as $\bm{u}^{\theta}_{t,r}(\bm{x}_{t})=F^{\theta}(\bm{x}_{t},t,r)$, and estimate both the one-step prediction and the two steps in the target with neural networks $F^{\theta}$, we thus obtain the $l(\bm{x}_{t},r,s,t;\theta)$ in SCD through Eq. 2. Due to equi-spacing time points as ts=sr=ht-s=s-r=h, it reads
@@
where the loss metric is set as squared $l_{2}$-distance, $w=\frac{1}{4h^{2}}$, and $\hat{\bm{x}}_{s}=\bm{x}_{t}+(s-t)\bm{u}^{\theta}_{s,t}(\bm{x}_{t})$, with details shown in Appendix B.2. Moreover, for IMMs, conditional samples $\{\bm{x}^{(i)}_{0},\bm{\varepsilon}^{(i)}\}_{i=1}^{B}$ within a mini-batch of size $B$ are first partitioned into different groups, and $\{r,s,t\}$ are drawn for each group. With similar flow map construction to CTs, the loss metric is implemented with a grouped kernel function as to minimize MMD (Gretton et al., 2012), applied to measure both inter- and intra-sample similarities between $\{\bm{\hat{x}}_{r},\bm{x}^{\theta}_{r}\}$ within the group, as further detailed in Appendix B.3.
@@ 2.4 Discussion: shortcutting flow paths under marginal velocity fields
However, such pairs $\{\bm{x}_{t},\bm{x}_{r}\}$ cannot be obtained via simulation-based sampling: once $\bm{x}_{t}$ is sampled, $\bm{x}_{r}$ remains inaccessible because both $\bm{v}_{t}(\bm{x})$ and its integral from $t$ to $r$ are intractable. To overcome this, a common design paradigm is employed, which is to let the network’s outputs, or the conditional velocity alternatively, estimate $\bm{x}_{r}$ in two steps: first producing an intermediate $\hat{\bm{x}}_{s}$, and then constructing an estimated target $\hat{\bm{x}}_{r}$, as shown in Eq. 5 in Sec. 2.2. This makes training feasible. Although one may also construct multi-step (i.e., more than two) flow map targets for simulating the $\{\bm{x}_{t},\bm{x}_{r}\}$ pairs (Kim et al., 2023), the paradigms of two-step target construction approximated by one-step prediction are sufficiently general according to the following theoretical justification with detailed proof in Appendix C.2. Note that we classify the aforementioned methods’ training objective into DTSC and CTSC. For example, $l_{\text{mf}}$ and $l_{\text{sct}}$ are instances of $l_{\text{ctsc}}$.
@@
From another perspective, this explains why distilling from a pretrained diffusion model is often more effective than training from scratch (Song et al., 2023; Lu and Song, 2025). Unlike (s)CT, which are trained from scratch, (s)CM benefits from distillation by learning from a pretrained velocity field $\bm{v}_{t}^{\phi}(\bm{x})$. In practical training, the conditional velocity $\bm{v}_{t|0}$ and network output $\bm{v}_{t}^{\theta}$ in $\mathrm{sg}(\cdot)$ in Eq. 6 and 9 are replaced with $\bm{v}_{t}^{\phi}$, which closely approximates $\bm{v}_{t}(\bm{x}_{t})$. This substantially reduces errors in estimating the two-step flow targets, providing more accurate supervision for network training.
@@ 3 Elucidating the design space of shortcut models
Empirically, we evaluate the proposed formulation using a unified codebase implementation with the same training iterations and batch sizes. For unconditional generation on CIFAR-10, we employ U-Nets (Song et al., 2020) ($\sim$55M param.) as the network architecture operating directly in the pixel space. For conditional generation in ImageNet-$\times256$, with and without classifier-free guidance, we use a SiT-B/2 (Ma et al., 2024) architecture ($\sim$131M param.), operating in the latent space via a pretrained VQVAE (Rombach et al., 2021). While sCT is originally initialized from the teacher diffusion model as stated in Lu and Song (2025), we train all the discussed models from scratch, for a fair comparison. Fig. 4(a) summarizes the results of one-step generation on the two datasets, with additional setting of classifier-guidance-free learning, as discussed in Geng et al. (2025a). Further details on settings are provided in Appendix D.1.
@@
Linear paths are generally regarded as more analytically tractable and easier to employ for training and sampling tricks (e.g., classifier-free guidance), owing to their simple formulation. By contrast, in pixel-space generative modeling, cosine paths are often considered more stable for training convergence, because they induce a stochastic process with fixed variance. Exploration of these two flow paths in the context of shortcut models remains underexplored. Here we extend cosine-path-based models (CT and sCT) to their linear-path counterparts. Fig. 4(a) shows that shortcut models with linear paths are more competitive. We attribute this to the fact that the marginal velocity fields generated by linear paths as conditional paths are at lower convex transport cost (Liu et al., 2022), implying lower curvature of the velocity-field-governed trajectories. Consequently, the simulated two-step flow map targets are less likely to deviate from the ideal. Furthermore, while cosine paths are optimal in the setting of diffusion and flow matching (Santos and Lin, 2023) under Fisher information metrics, we theoretically justify that linear paths in the setting of shortcut models are optimal conditioned on data samples in Appendix C.3. Based on this, our subsequent analysis will focus on the linear path.
@@
Under mild regularity conditions shown in Appendix C.4.1, the Wasserstein-2 distance of shortcut models with one-step generation is bounded as: $W_{2}^{2}(p_{0},p^{\theta}_{0})\leq 2\left(\mathrm{BV_{\text{ctsc}}}+8\mathrm{Var}\big[\frac{d}{dt}\bm{u}_{t,r}^{\theta}(\bm{x}_{t})\big]+8\sigma_{\bm{v}_{t|0}}^{2}\right)\Big|_{r=0,t=1},$ (10) $W_{2}^{2}(p_{0},p^{\theta}_{0})\leq 2\Big(\mathrm{BV_{\text{dtsc}}}+8\delta_{2}^{2}\,\mathrm{Var}\big[\bm{u}_{s,r}^{\theta}(\bm{x}_{t})\big]+8(1+\ell^{2}\delta_{2}^{2})\delta_{1}^{2}\,\sigma_{\text{dtsc}}^{2}\Big)\Big|_{r=0,t=1},$~ (11) where $\mathrm{BV_{\bullet}}=\mathrm{Bias}^{2}_{\bullet\text{-tgt}}+\mathrm{Bias}^{2}_{\bullet\text{-loss}}+2\mathrm{Var}\!\big[\bm{u}_{t,r}^{\theta}(\bm{x}_{1})\big]$ with $\bullet\in\{\mathrm{ctsc},\mathrm{dtsc}\}$, and $\mathrm{Bias}^{2}_{\bullet\text{-tgt}}$ and $\mathrm{Bias}^{2}_{\bullet\text{-loss}}$ are defined in Prop.C.8 ; $\delta_{1}=t-s$, $\delta_{2}=s-r$; $\ell$ is the local Lipschitz constant of $\bm{u}^{\theta}$; $\sigma_{\bm{v}_{t|0}}^{2}$ is the variance of the conditional velocity, defined by $\sigma_{\bm{v}_{t|0}}^{2}\coloneqq\mathrm{Var}(\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0}))$; $\sigma_{\text{dtsc}}^{2}=\sigma_{\bm{v}_{t|0}}^{2}$ for CT’s two-step flow map targets, or $\sigma_{\text{dtsc}}^{2}=\mathrm{Var}\!\big[\bm{u}_{t,s}^{\theta}(\bm{x}_{t})\big]$ when using SCD’s flow map targets.